# Feature Alignment: Rethinking Efficient Active Learning via Proxy in the Context of Pre-trained Models

### v1

@@ 5 Method
In line with the analysis presented in sec. 4, addressing the performance drop of SVPp primarily involves focusing on the disparities between SVPp and standard active learning in redundant and missing regions. For samples in the redundant region, the fine-tuned model can predict these samples correctly with high confidence, but the proxy model struggles to differentiate them. This suggests that the features of the fine-tuned model are more suitable for distinguishing these samples. So, we suggest performing a complete model fine-tuning as the number of these samples increases, updating the pre-computed features to enhance the proxy model’s discriminative capacity for this region and, consequently, reducing the selection of redundant region samples. The data selection stage of our method is shown in fig. 6.
Figure 6: The Data Selection Pipeline in our approach, Aligned Selection via Proxy (ASVP). After acquiring new labels from the oracle, we assess the necessity of updating the pre-computed features. If required, the pre-trained model is fine-tuned, and the pre-computed features are subsequently updated.
For the missing region, these samples are selected to assist the fine-tuned model in preserving valuable pre-trained information as discussed in sec. 4. So, the focus should be on aiding the fine-tuned model in inheriting the ability to distinguish missing region samples from the pre-trained model rather than selecting them. Specifically, in sec. 5.1, we describe how to determine the moment for updating pre-computed features. In sec. 5.2, we detail how to mitigate the impact of the missing samples by altering the training method of the final model.
5.1 Aligning Proxy Features
As discussed in sec. 4, with the increase in annotations, the quality of fine-tuned features improves gradually, resulting in a higher proportion of samples selected by the proxy model from the redundant region. Also, adding the samples from the redundant region rarely improves the performance of the fine-tuned model. Hence, we assess the ratio of redundant region samples by evaluating the changes in the performance of the fine-tuned model after adding more annotated samples. To assess the changes in the fine-tuned model’s performance, we employ LogME-PED [41, 22] as an indicator. LogME-PED is an evaluation metric for pre-trained models, where the LogME component estimates the pre-trained model’s capability to effectively differentiate given annotated samples, and PED, a physically inspired model, simulates feature learning dynamics through multiple iterations.
@@ 5.1 Aligning Proxy Features
5.2 Preserving Valuable Pre-trained Features
As speculated in sec. 4, sampling from the missing region improves model performance by reducing the alterations of the pre-trained model by fine-tuning. Based on this, we propose preserving valuable and sensitive pre-trained information by adjusting the training method of the final model, instead of trying to incorporate these samples. Specifically, after sample selection, we employ the LP-FT method [20] to train the final model. LP-FT initiates the training process by exclusively updating the linear head with frozen pre-trained weights, followed by a standard fine-tuning phase. This sequential approach is designed to minimize alterations to the pre-trained model induced by fine-tuning.
Furthermore, the missing region comprises two types of samples: those the proxy model based on pre-trained features can correctly predict (region B2) and those it cannot (region B1). Similar to the redundant region analysis, as annotations increase, fine-tuned features outperform pre-trained ones. This leads to a gradual increase in samples from region B1 compared to region B2. In this scenario, continuing to use LP-FT may retain incorrect pre-trained information. Therefore, we also switch the model’s training method from LP-FT to FT, while updating pre-computed features.
@@ 6 Results
Our method is validated on four datasets: ImageNet-1k [11], CIFAR-10 [19], CIFAR-100 [19] and Oxford-IIIT Pet dataset [27] with four typical active learning strategies, The experiment setup is clarified in sec. 6.1. We propose the new efficient active learning evaluation metric in sec. 6.2. The results are shown in sec. 6.3 and sec. 6.4. The ablation study is shown in sec. 6.5. The detailed study of the impact of the position of updating features is shown in sec. 6.6. Additionally, given the efficiency of our framework, decreasing the amount of sample selected in each active learning iteration to alleviate the redundant problem of batch active learning strategy is feasible. The results are shown in sec. 6.7.
@@ 6.1 Experiment Setup
Active learning strategies. To evaluate the compatibility of our proposed ASVP with existing active learning strategies, four typical active learning strategies are included: (1) Margin: uncertainty-based sampling, one of the SOTA in the context of fine-tuning pre-trained models [30], (2) Confidence: uncertainty-based sampling, (3) BADGE: combination of uncertainty and diversity [1], (4) ActiveFT(al), feature space coverage with features from fine-tuned model [37].
Baseline. (1) Standard active learning: training the whole model with fine-tuning (FT) or linear-probing then fine-tuning (LP-FT) [20] in each active iteration, (2) SVPp [42]: selecting samples via MLP proxy model based on pre-trained features and training the final model with FT, (3) ActiveFT(pre) [37]: a single-shot active learning strategy that selects all sample in single iteration based on pre-trained features.
Implementations. We utilized the ResNet-50 model [16], pre-trained on ImageNet using the BYOL-eman [4] method in all our experiments. The training hyperparameters of the two model training approaches, fine-tune (FT) and LP-FT, are provided in appendix 8. For the baseline method (SVPp) and our method (ASVP), we employed 2-layer MLPs as proxy models, where its architecture is Linear + BatchNorm + ReLU + Linear. The hidden layer width matched the input feature dimensions. For the ImageNet, CIFAR-10, and Pets datasets, we conducted 10 active learning iterations, with each iteration selecting 10,000, 200, and 200 samples, respectively. In the CIFAR-100, 15 active learning iterations were performed, with each iteration selecting 400 samples. The difference in LogME-PED scores between consecutive active learning iterations is set as 1 for the threshold of updating pre-computed features.
@@ 6.2 Metric
To this end, we propose a new evaluation metric: equivalent saving amount. Let $N_{1}$ represent the number of samples selected by the active learning strategy, and let $A$ be the accuracy achieved by a model trained with these $N_{1}$ labeled samples. We utilize interpolation to estimate the number of samples, $N_{2}$, required to achieve accuracy $A$ for a random baseline. The number of saved samples is $N_{2}-N_{1}$ and sample saving ratio is $\frac{N_{2}-N_{1}}{N_{2}}$. The overall cost of active learning, including annotation and training costs, equals $N_{1}\times P_{l}+C_{tr}$, where $P_{l}$ denotes the unit price of annotation and $C_{tr}$ represents the training cost of active learning. Additionally, in real-world scenarios, active learning involves multiple iterations with varying counts. We calculate the average sample saving ratio to assess the performance.
In this paper, we primarily present the average sample savings ratio, overall costs, and computational time. Additionally, detailed accuracy and the corresponding savings in sample count in each active learning iteration are provided in appendix 9.
@@ 6.3 Sample Saving Ratio and Overall Cost
The average sample savings ratios are presented in table 1, with each experimental setting repeated three times to report both the average and standard deviation. The table also illustrates the total cost required for active learning to reach the accuracy achieved by the random baseline at the last active learning iteration (i.e., the accuracy achieved when randomly selecting 100k, 2000, 6000, and 2000 samples from ImageNet, CIFAR-10, CIFAR-100, and Pets datasets, respectively). The training cost is assessed based on AWS EC2 P3 instances [42], while the annotation cost is evaluated using AWS Mechanical Turk, with three annotations per sample to ensure labeling quality. Additionally, the accuracy and corresponding saving in label counts on ImageNet with margin sampling are shown in fig. 7. The other results are shown in appendix 9.
Selection Training AL ImageNet CIFAR-10 CIFAR-100 Pets Cost
Method Method Strategy Avg. Sample Cost Avg. Sample Cost Avg. Sample Cost Avg. Sample Cost Saving
Saving Ratio Saving Ratio Saving Ratio Saving Ratio Ratio
Standard FT Random - 3600 - 72 - 216 - 72 100%
Standard FT Margin 22.70$\pm$1.90 2739 42.77$\pm$1.05 44 8.86$\pm$2.25 207 31.67$\pm$2.30 48 75%
BADGE 16.87$\pm$0.67 2962 35.95$\pm$0.78 46 8.97$\pm$2.54 212 33.01$\pm$3.17 47 77%
Confidence -19.82$\pm$1.05 3672 20.09$\pm$3.06 54 -5.24$\pm$3.16 226 27.24$\pm$3.28 50 88%
ActiveFT(al) - - 9.32$\pm$3.43 62 3.09$\pm$0.44 379 -5.20$\pm$2.28 141 152%
Standard LP-FT Margin 30.97$\pm$0.51 2461 38.83$\pm$0.45 45 17.31$\pm$1.90 203 6.97$\pm$2.55 74 82%
BADGE 24.23$\pm$1.11 2777 25.35$\pm$1.32 51 15.35$\pm$1.55 209 6.74$\pm$3.84 70 85%
Confidence -12.71$\pm$2.66 3636 14.69$\pm$4.65 52 4.88$\pm$4.22 229 9.58$\pm$3.41 73 95%
ActiveFT(al) - - 5.38$\pm$1.81 73 8.08$\pm$0.64 478 -29.33$\pm$10.58 148 176%
Single-shot FT ActiveFT(pre) - - -3.36$\pm$4.92 75 1.89$\pm$1.68 434 -11.92$\pm$1.26 167 170%
LP-FT - - 19.58$\pm$2.21 189 10.86$\pm$1.24 410 -38.10$\pm$2.61 265 273%
SVPp FT Margin 16.64$\pm$2.06 2795 23.63$\pm$4.01 53 1.36$\pm$1.45 213 23.48$\pm$4.09 50 80%
BADGE 13.01$\pm$1.36 3038 14.96$\pm$5.89 58 -2.11$\pm$0.85 376 -7.78$\pm$2.42 129 130%
Confidence -25.97$\pm$1.28 6756 -11.54$\pm$2.30 73 -14.43$\pm$2.49 496 2.40$\pm$7.25 67 153%
ActiveFT(al) - - -4.72$\pm$5.78 72 0.25$\pm$0.92 388 -11.63$\pm$6.85 519 333%
ASVP FT/LP-FT Margin 30.22$\pm$1.85 2381 35.03$\pm$1.89 49 13.31$\pm$1.86 200 21.98$\pm$0.50 49 74%
BADGE 24.43$\pm$1.12 2727 33.67$\pm$0.84 47 8.16$\pm$1.98 208 6.10$\pm$5.59 57 79%
Confidence -6.37$\pm$1.23 3326 23.92$\pm$5.16 52 -0.62$\pm$1.43 362 14.81$\pm$3.17 51 100%
ActiveFT(al) - - 27.45$\pm$1.42 53 11.28$\pm$2.01 217 -7.74$\pm$4.81 76 93%
@@
(a) Accuracy
(b) Sample Savings
Figure 7: The accuracy and equivalent non-active learning label amounts of the standard method, SVPp and our method ASVP on the ImageNet. The equivalent non-active learning label amount refers to the number of samples selected by a random baseline to achieve the same accuracy as active learning.
@@ 6.4 Computational Time
We define the average model training time for both the fine-tuned model and the proxy model in each active learning iteration as $T_{tr,full}$ and $T_{tr,proxy}$, respectively. The time for processing all unlabeled samples through the whole pre-trained model and the proxy model to compute active learning metrics and select samples is denoted as $T_{f,full}$ and $T_{f,proxy}$, respectively. The time spent on forwarding all samples through the model to record features is represented by $T_{pre}$. The active learning iteration count is given by $N_{al}$. The total active learning time for standard active learning, $T_{s,full}$, is $N_{al}\times(T_{tr,full}+T_{f,full})$. The total time for SVPp, $T_{s,svpp}$, is $N_{al}\times(T_{tr,proxy}+T_{f,proxy})+T_{pre}$. For ASVP, $T_{s,asvp}$, the total time is $N_{al}\times(T_{tr,proxy}+T_{f,proxy})+2\times T_{pre}+T_{tr,full}$. When maintaining the same number of active learning iterations as standard active learning, the time saved by ASVP is primarily influenced by $2\times T_{pre}+T_{tr,full}$.
$T_{pre}$ $N_{al}\times T_{f}$ $N_{al}\times T_{tr}$ $T_{s}$
$/hrs$ $/hrs$ $/hrs$ $/hrs$
Standard - 11.66 23.86 35.52
SVPp 0.92 0.07 0.11 1.10
ASVP 1.84 0.07 2.67 4.58
@@
Given that the computation time of the proxy model, MLP, is significantly less than that of the fine-tuned model, ASVP also opens up the potential for increasing the number of active learning iterations. The computational times for standard active learning, SVPp, and ASVP are presented in table 2, where the running time is measured on ImageNet. Since our method achieves the same accuracy as the standard pipeline by only adopting LP-FT, the running time of ASVP is not reported based on practical running. This time is estimated based on the SVPp running time. The computational time on other datasets is shown in appendix 10.

### v2

@@ 5 Method
This section first addresses the issue of the increasing proportion of region C samples, which makes LP-FT insufficient to prevent the decline in SVPp active learning performance. Region C refers to samples that the fine-tuned model can predict correctly with high confidence, while the proxy model struggles to differentiate them (high uncertainty). This indicates that the fine-tuned model’s features are more suitable for distinguishing these samples. To address this, we recommend fine-tuning the pre-trained model as the number of region C samples increases. Updating the pre-computed features enhances the proxy model’s discriminative capacity, thereby reducing the proportion of region C samples. Additionally, when updating pre-computed features, we switch the model training method from LP-FT to fine-tuning. Continuing with LP-FT when fine-tuned features outperform pre-trained features can lead to suboptimal results by hindering necessary modifications to the pre-trained model.
The ASVP framework is divided into three stages: pre-computation of features, data selection, and final model training. The data selection stage of our method is illustrated in fig. 10. During the data selection stage, the proxy model uses pre-computed features (pre-computed features refer to pre-trained features when feature updating is not triggered) as input and performs multiple rounds of active learning (alternating between sample selection and proxy model training). At each iteration, we check if an update to the pre-computed features is needed. If the annotation budget is exhausted without triggering a feature update, LP-FT is used to train the final model. If an update is detected, fine-tuning is performed to generate new pre-computed features for the following rounds of sample selection. After the sample selection, the final model is trained using fine-tuning.
Figure 10: The Data Selection Pipeline in our approach, Aligned Selection via Proxy (ASVP). After acquiring new labels from the oracle, we assess the necessity of updating the pre-computed features. If required, the pre-trained model is fine-tuned, and the pre-computed features are updated.
Next, we explain when it becomes necessary to update the pre-computed features. As discussed in sec. 4, as the number of annotations increases, the quality of fine-tuned features improves gradually, resulting in a higher proportion of samples selected by the proxy model from the redundant region. Also, including the samples from the redundant region rarely improves the performance of the fine-tuned model. Hence, we assess the ratio of redundant region samples by evaluating the changes in the performance of the fine-tuned model after adding more annotated samples. To assess the changes in the fine-tuned model’s performance, we employ LogME-PED (You et al., 2021; Li et al., 2023) as an indicator.
@@ 6 Results
Our method is validated on four datasets: ImageNet-1k (Deng et al., 2009), CIFAR-10 (Krizhevsky et al., 2009), CIFAR-100 (Krizhevsky et al., 2009) and Oxford-IIIT Pet dataset (Parkhi et al., 2012) with five typical active learning strategies. The experiment setup is clarified in sec. 6.1. We propose the new efficient active learning evaluation metric in sec. 6.2. The results are shown in sec. 6.3 and sec. 6.4. The ablation study is shown in sec. 6.5. The detailed study of the impact of the position of updating features is shown in sec. 6.6. Additionally, given the efficiency of our framework, decreasing the amount of sample selected in each active learning iteration to alleviate the redundant problem of batch active learning strategy is feasible. The results are shown in appendix F.
@@ 6.1 Experiment Setup
Active learning strategies. To evaluate the compatibility of our proposed ASVP with existing active learning strategies, four typical active learning strategies are included: (1) Margin: uncertainty-based sampling, one of the SOTA in the context of fine-tuning pre-trained models (Scheffer et al., 2001), (2) Confidence: uncertainty-based sampling (Wang & Shang, 2014), (3) Coreset: diversity-based sampling (Sener & Savarese, 2018), (4) BADGE: combination of uncertainty and diversity (Ash et al., 2020), (5) ActiveFT(al), feature space coverage with features from fine-tuned model (Xie et al., 2023).
Baseline. (1) Standard active learning: training the whole model with fine-tuning (FT) or linear-probing then fine-tuning (LP-FT) (Kumar et al., 2022) in each active iteration, (2) SVPp (Zhang et al., 2024): selecting samples via MLP proxy model based on pre-trained features and training the final model with FT, (3) ActiveFT(pre) (Xie et al., 2023): a single-shot active learning strategy that selects all sample in single iteration based on pre-trained features.
Implementations. We utilized the ResNet-50 model (He et al., 2016), pre-trained on ImageNet using the BYOL-EMAN (Cai et al., 2021) method in all our experiments. The training hyper-parameters of the two model training approaches, fine-tune (FT) and LP-FT, are provided in appendix A. For the baseline method (SVPp) and our method (ASVP), we employed 2-layer MLPs as proxy models, where its architecture is Linear + BatchNorm + ReLU + Linear. The hidden layer width matched the input feature dimensions. We performed multiple active learning iterations until the model performance approached that of training on the entire dataset (upper bound). Specifically, for the ImageNet, CIFAR-10, CIFAR-100, and Pets datasets, we conducted 16, 18, 22, and 10 active learning iterations, respectively. The query step was set to 200 samples per iteration for the Pets dataset. For ImageNet, CIFAR-10, and CIFAR-100, we used varying query steps: for ImageNet, we queried 10,000 samples per iteration until reaching 100,000 samples, after which we queried 50,000 samples per iteration. For CIFAR-10, we queried 200 samples per iteration until reaching 2,000 samples, then increased to 500 samples per iteration. For CIFAR-100, we queried 400 samples per iteration until reaching 6,000 labeled samples, then switched to 2,000 samples per iteration. The difference in LogME-PED scores between consecutive active learning iterations is set as 1 for the threshold of updating pre-computed features.
@@ 6.2 Metric
To this end, we propose a new evaluation metric: equivalent saving amount. Let $N_{1}$ represent the number of samples selected by the active learning strategy, and let $A$ be the accuracy achieved by a model trained with these $N_{1}$ labeled samples. We utilize interpolation to estimate the number of samples, $N_{2}$, required to achieve accuracy $A$ for a random baseline. We refer to $N_{2}$ as the equivalent number of non-AL labels in the subsequent discussions. The number of saved samples is $N_{2}-N_{1}$ and sample saving ratio is $\frac{N_{2}-N_{1}}{N_{2}}$. Subsequently, we compute the average sample saving ratio across all iterations of active learning to assess the performance of the active learning strategy. Furthermore, the overall cost of active learning, including annotation and training costs, equals $N_{1}\times P_{l}+C_{tr}$, where $P_{l}$ denotes the unit price of annotation and $C_{tr}$ represents the training cost of active learning.
@@ 6.3 Sample Saving Ratio and Overall Cost
The average sample savings ratios are presented in table 1, with each experimental setting repeated three times to report both the average and standard deviation. The table also illustrates the total cost required for active learning to reach the accuracy achieved by the random baseline at the last active learning iteration (i.e., the accuracy achieved when randomly selecting 400k, 6000, 20000, and 2000 samples from ImageNet, CIFAR-10, CIFAR-100, and Pets datasets, respectively). The training cost is assessed based on AWS EC2 P3 instances (Zhang et al., 2024), while the annotation cost is evaluated using AWS Mechanical Turk, with three annotations per sample to ensure labeling quality. Additionally, Figures 11 and 12 show the accuracy and corresponding label savings achieved with margin sampling, while results using other active learning strategies are provided in Appendix E. Figure 13 shows the sampling time and total cost required by standard active learning, SVPp, ASVP (ours), and the single-shot active learning strategy, ActiveFT (pre), to achieve the accuracy attained by the random baseline at the final active learning iteration.
@@
ImageNet CIFAR-10 CIFAR-100 Pets Cost
Selection Training AL Avg. Sample Cost Avg. Sample Cost Avg. Sample Cost Avg. Sample Cost Saving
Method Method Strategy Saving Ratio (\$) Saving Ratio (\$) Saving Ratio (\$) Saving Ratio (\$) Ratio
Standard FT Random 0% 14400 0% 216 0% 720 0% 72 100%
Standard FT Margin 30.30$\pm$1.09 8752 40.23$\pm$0.51 134 14.54$\pm$2.46 578 31.67$\pm$2.30 49 68%
BADGE 21.33$\pm$0.38 10018 36.18$\pm$0.34 128 13.73$\pm$2.07 580 33.01$\pm$3.17 47 69%
Confidence -3.42$\pm$0.75 9619 24.99$\pm$1.37 142 4.23$\pm$2.34 567 27.24$\pm$3.28 50 70%
Coreset - - 12.44$\pm$2.42 192 -13.90$\pm$2.31 722 -4.05$\pm$2.31 74 97%
ActiveFT(al) - - 12.39$\pm$3.26 186 3.03$\pm$1.17 798 -5.20$\pm$2.28 97 111%
Standard LP-FT Margin 32.22$\pm$0.27 9884 37.39$\pm$1.28 139 22.31$\pm$1.83 509 6.97$\pm$2.55 73 76%
BADGE 24.12$\pm$0.54 11901 29.54$\pm$0.70 136 19.81$\pm$1.28 528 6.74$\pm$3.84 68 78%
Confidence 0.11$\pm$1.51 10125 20.10$\pm$4.01 138 13.49$\pm$2.90 510 9.58$\pm$3.41 73 77%
Coreset - - -4.60$\pm$4.09 244 -16.76$\pm$1.78 769 -36.93$\pm$4.53 110 124%
ActiveFT(al) - - -2.13$\pm$0.11 249 6.38$\pm$1.35 755 -29.33$\pm$10.58 104 122%
Single-shot FT ActiveFT(pre) - - 1.40$\pm$1.50 235 -0.21$\pm$1.41 822 -11.92$\pm$1.26 118 129%
LP-FT - - 5.02$\pm$4.02 238 8.92$\pm$2.34 742 -38.10$\pm$2.61 235 180%
SVPp FT Margin 25.16$\pm$1.47 9094 19.86$\pm$2.53 184 3.51$\pm$0.77 654 23.48$\pm$4.09 50 77%
BADGE 16.17$\pm$1.84 11374 19.35$\pm$3.66 156 0.51$\pm$0.74 657 -7.78$\pm$2.42 85 90%
Confidence -8.56$\pm$2.88 9678 1.53$\pm$1.82 171 -7.93$\pm$1.85 644 2.40$\pm$7.25 66 82%
Coreset - - -54.78$\pm$8.88 1355 -56.53$\pm$3.12 903 -38.63$\pm$11.64 87 291%
ActiveFT(al) - - 3.79$\pm$4.31 192 -0.50$\pm$1.42 713 -11.63$\pm$6.85 319 210%
ASVP FT/LP-FT Margin 34.38$\pm$1.57 8690 32.45$\pm$1.13 127 12.82$\pm$1.40 590 21.98$\pm$0.50 47 67%
BADGE 25.57$\pm$2.54 10305 32.87$\pm$0.84 137 10.11$\pm$1.13 577 6.10$\pm$5.59 55 73%
Confidence 5.88$\pm$0.66 9479 25.67$\pm$2.48 167 2.75$\pm$1.17 616 14.81$\pm$3.17 50 75%
Coreset - - 11.11$\pm$1.74 192 -26.16$\pm$1.76 767 -46.3$\pm$10.26 91 107%
ActiveFT(al) - - 26.67$\pm$0.66 155 6.81$\pm$1.70 775 -7.74$\pm$4.81 47 82%
(a) ImageNet
(b) CIFAR-10
(c) CIFAR-100
(d) Pets
Figure 11: The accuracy of the standard method, SVPp, and our method ASVP on (a) ImageNet, (b) CIFAR-10, (C) CIFAR-100 and (d) Pets using margin sampling.
(a) ImageNet
(b) CIFAR-10
(c) CIFAR-100
(d) Pets
Figure 12: The equivalent number of non-active learning label amounts comparison of the standard method, SVPp, and our method ASVP on (a) ImageNet, (b) CIFAR-10, (C) CIFAR-100 and (d) Pets using margin sampling.
(a) ImageNet
(b) CIFAR-10
(c) CIFAR-100
(d) Pets
Figure 13: Computation efficiency and overall cost comparison of the standard active learning method, SVPp, and our method ASVP on (a) ImageNet, (b) CIFAR-10, (C) CIFAR-100 and (d) Pets using margin sampling.
@@ 6.4 Computational Time
We define the average training time for the fine-tuned and the proxy models in each active learning iteration as $T_{tr,full}$ and $T_{tr,proxy}$, respectively. The time for processing all unlabeled samples through the whole pre-trained model and the proxy model to compute active learning metrics and select samples is denoted as $T_{f,full}$ and $T_{f,proxy}$, respectively. The time spent on forwarding all samples through the model to record features is represented by $T_{pre}$. The number of active learning iteration is given by $N_{al}$. The total active learning time for standard active learning, $T_{s,full}$, is $N_{al}\times(T_{tr,full}+T_{f,full})$. The total time for SVPp, $T_{s,svpp}$, is $N_{al}\times(T_{tr,proxy}+T_{f,proxy})+T_{pre}$. For ASVP, the initial pre-computed features are obtained from a pre-trained model. When updating these pre-computed features is necessary, the model is fine-tuned on labeled data, and the pre-computed features are refreshed accordingly. Let $N_{pre}$ denote the number of times the ASVP features are pre-computed, then, the time required for pre-computing features is $N_{pre}\times T_{pre}+(N_{pre}-1)\times T_{tr,full}$. The total time required by ASVP, denoted as $T_{s,asvp}$, is $N_{al}\times(T_{tr,proxy}+T_{f,proxy})+N_{pre}\times T_{pre}+(N_{pre}-1)\times T_{tr,full}$.
@@
$T_{pre}(hrs)$ $N_{al}\times T_{f}(hrs)$ $N_{al}\times T_{tr}(hrs)$ $T_{s}(hrs)$
Standard - 16.36 80.97 97.33
SVPp 0.92 0.10 0.46 1.48
ASVP 8.60 0.10 0.46 9.16
The computation time of standard active learning, SVPp, and ASVP on ImageNet is presented in table 2. The results on other datasets are shown in appendix B. The time for ASVP mainly arises from updating pre-computed features, as the computation time for the proxy model, an MLP, is significantly lower than that of the fine-tuned model. This means that increasing the number of active learning iterations $N_{al}$ has a marginal impact on ASVP’s total time. Therefore, this opens up the potential to improve the diversity of selected samples by reducing the number of samples chosen in each active learning iteration.